HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

INVESTIGATION OF THE MICROSOLVATION OF SECONDARY ORGANIC AEROSOL PRECUSORS: MICROWAVE SPECTROSCOPY OF

NITROMETHOXYPHENOLS HYDRATES

\underline{C. BRACQUART} , M. CHRAYTEH, A. CUISSET, Université du Littoral Côte

d´ Opale, UR4493, LPCA, Laboratoire de Physico-Chimie de l´ Atmosphère, F-59140 Dunkerque, France; \underline{J. A. CLAUS}, A. DEGUINE, M. GOUBET, Univ. Lille, CNRS, UMR8523 - PhLAM - Physique des Lasers Atomes et Molécules,

F-59000 Lille, France

Atmospheric aerosol particles strongly influence the Earth´ s atmosphere, and their contribution to climate change is versatile. Especially, Secondary Organic Aerosols (SOAs) play a major role in the radiative budget and the clouds formation through their hygroscopic properties and atmospheric chemistry.^{1} Methoxyphenols (lignin´ s monomer) represent a significant part of biomass burning emissions. Recent works have shown that SOAs are formed by oxidation of methoxyphenols such as guaiacol and nitro-guaiacols (NG) are the main oxydation products yielding to SOAs with an hydrophobic character.^{2}

We present here the rotational spectroscopy of 4- and 5-NG isomers studied using a free jet Fabry-Perot Fourier-transform microwave (FP-FTMW) spectrometer (2 to 20 GHz) with the support of quantum chemistry calculations. The observation of the hyperfine structure induced by the presence of a ^{14}N nucleus allowed us to simulate the nuclear quadrupole coupling constants in addition to the Watsonian rotationnal parameters. Each isomer shows only one planar conformer stabilized by an intramolecular hydrogen bond as already observed for guaiacol.^{3} However, the position of the nitro-group seems to influence the global stability of the molecule. The microsolvation of 4-NG with water has been investigated, unveiling only one monohydrate complex in agreement with the demonstrated hydrophobic character of 4-NG aerosols.^{3} The study of the hydration of 5-NG is ongoing as well as the study of the less volatile 4-nitro-syringol, a precursor with more hydrophilic SOAs than NG.^{4}

1doi:10.1016/B978-0-12-257060-5.X5000-X, Barbara J. Finlayson-Pitts and James N. Pitts, Jr., Chemistry of the Upper and Lower Atmosphere, Academic, 951 (2000)

2doi:10.1364/AO.56.00E116, W. Ahmad et al., Infrared spectroscopy of secondary organic aerosol precursors and investigation of the hygroscopicity of SOA formed from the OH reaction with guaiacol and syringol, Applied Optics, 56, E116-E122 (2017)

3doi:10.1063/1.5089426, A. Jabri et al., Conformational landscape and inertial defect of methoxyphenol isomers studied by mm-wave spectroscopy and quantum chemistry calculations, The Journal of Chemical Physics, 150,10, 104303 (2019)

4This work is supported by the LabEx CaPPA (11-LABX-0005-01) and by the CPER ECRIN.

p-number: p159  Submitted on Thu, 15 Jun 2023 11:32:48 +0200